OpenStudy (anonymous):

Find sin x/2, cos x/2, and tan x/2 from the given information. cos x = −15/17, 180° < x < 270° sin x/2= cos x/2= tan x/2=

OpenStudy (anonymous):

\[2 \cos ^{2}\frac{ x }{ 2 }=1+\cos x=1-\frac{ 15 }{ 17 }=\frac{ 2 }{17 },\cos^{2}\frac{ x }{ 2 } =\frac{ 1 }{ 17 }\] \[\cos \frac{ x }{ 2 }=-\frac{ 1 }{ \sqrt{17} } ....(1)\] \[\sin \frac{ x }{ 2 }=\frac{ 1 }{ \sqrt{17} },90<\frac{ x }{ 2 }<135\] \[\sin \frac{ x }{ 2 }>0,\cos \frac{ x }{ 2 }<0,\tan \frac{ x }{ 2 }<0\] \[2\sin^{2}\frac{ x }{ 2 }=1-\cos x=1-\frac{ -15 }{ 17 }=\frac{ 32 }{ 17 }, \sin \frac{ x }{ 2 }=\frac{ 4 }{\sqrt{17} } .....(2)\] divide (2) by (1) and get the value of tan x/2